For Two Natural Numbers x and y, if 2x-1 <= 4 and 2x+2y <= 7. Find the possible values of y.
Answer by fcabanski(1391)

Natural numbers are the counting numbers: 1, 2, 3....


2x-1 <= 4


2x <= 5


x<=5/2 so x is 2 or 1.


2x+2y <= 7(use x=1)


2 + 2y <= 7


2y <= 5


y <= 5/2 so y = 2 or 1.


2x+2y <= 7 (use x=2)


4 + 2y <= 7


2y <= 3


y <= 3/2 so y = 1


Sometimes 0 is considered a natural number. In that case, any of the above could include 0 for y. And if x = 0:


2y <= 7


y <= 7/2 so y could be 3, 2, 1 or 0